# A Bayesian Dirichlet Auto-Regressive Conditional Heteroskedasticity Model for Compositional Time Series

### v1

@@ 5 Discussion
We introduced the Bayesian Dirichlet Auto-Regressive Moving Average with Dirichlet Auto-Regressive Conditional Heteroskedasticity (B-DARMA-DARCH) model, a novel approach for modeling and forecasting dynamic compositional time series with time-varying volatility. By incorporating a DARCH component into the B-DARMA framework, the model effectively captures both mean dynamics and heteroskedasticity inherent in compositional data.
Simulation studies demonstrated the superior performance of the B-DARMA-DARCH model across six scenarios, including data generated from DARMA, DARCH, and tVARMA processes with random shocks simulating misreported observations and regime changes. The B-DARMA-DARCH model consistently achieved the lowest forecast error metrics and exhibited minimal residual autocorrelation, indicating its robustness in capturing complex volatility dynamics, especially in the presence of heteroskedasticity.
Applying the B-DARMA-DARCH model to Airbnb’s currency fee proportions across different regions, we showed that it effectively captures temporal dynamics and volatility patterns in real-world data. The model generally outperformed the standard B-DARMA and B-tVARMA models in terms of forecast accuracy and residual diagnostics. Notably, the inclusion of the DARCH component allowed the model to account for significant disruptions, such as those caused by the COVID-19 pandemic, highlighting its robustness in the face of structural breaks and extreme events.
While the B-DARMA-DARCH model introduces additional complexity due to the dynamic modeling of the precision parameter, this complexity is justified by the improved model fit, empirical coverage rates, and forecasting accuracy in the presence of heteroskedasticity. The model offers a comprehensive framework that accounts for both mean dynamics and volatility patterns, which is particularly important in financial applications where volatility significantly influences decision-making and risk assessment.
However, the computational demands are higher, necessitating careful tuning of sampling parameters to ensure convergence and efficiency. Future research could focus on optimizing computational strategies or exploring approximate inference methods to mitigate these demands.
Several limitations warrant consideration. The current model models each region independently, even though some currencies are shared across multiple regions. This assumption of independence may not hold in practice, as economic events affecting a shared currency could simultaneously impact multiple regions. Incorporating hierarchical structures or accounting for inter-regional dependencies could enhance model performance by capturing these cross-regional correlations, albeit at the cost of increased computational complexity. Additionally, extending the model to handle zero-inflated compositional data or integrating external covariates that influence both the mean and volatility components could broaden its applicability.
@@ Acknowledgement
The authors thank Sean Wilson, Jess Needleman, Jenny Cheng, Erica Savage, Liz Medina, and Yuling Kuo for helpful discussions, Ellie Mertz, Adam Liss, and Peter Coles for championing the research, and Lauren Mackevich for their indispensable operational support.
Supplementary Materials
Alternative Log-Ratio Transformations
Although we have used the additive log-ratio (ALR) link in equations (3), (4), and (5), other transformations to and from the simplex may be employed. A common choice is the centered log-ratio (clr), defined as
${\textnormal{clr}}(\mathbf{y})\;=\;\Bigl{(}\,\ln\bigl{(}\tfrac{y_{1}}{g(\mathbf{y})}\bigr{)},\;\ln\bigl{(}\tfrac{y_{2}}{g(\mathbf{y})}\bigr{)},\;\dots,\;\ln\bigl{(}\tfrac{y_{J}}{g(\mathbf{y})}\bigr{)}\Bigr{)},$
where $g(\mathbf{y})=(\,y_{1}\,y_{2}\,\cdots\,y_{J})^{1/J}$ is the geometric mean of $\mathbf{y}$. The clr mapping places all $J$ components in a transformed $\mathbb{R}^{J}$ space but introduces the constraint that their sum is zero.
Another option is to use the Isometric Log-Ratio (ilr) transformation. For each index $j$, the ilr-transformed coordinate can be expressed as
$z_{j}=\sqrt{\frac{r_{j}}{r_{j}+1}}\,\ln\biggl{(}\frac{g_{j}(\mathbf{y)}}{\bigl{(}\prod_{i\in H_{j}}y_{i}\bigr{)}^{1/r_{j}}}\biggr{)},$
where $g_{j}(\mathbf{y})$ is defined as the geometric mean of a chosen subset $S_{j}\subseteq\mathbf{y}$, and $H_{j}$ is the complement of $S_{j}$ in $\mathbf{y}$. The value $r_{j}$ is the number of elements in $H_{j}$. The particular way in which $S_{j}$ and $H_{j}$ are selected depends on the specific application and how one wishes to interpret the ilr coordinates.
These mappings are mutually related by linear transformations (Egozcue et al., 2003) and each provides a bijection from the $(J-1)$-dimensional simplex onto a $(J-1)$-dimensional subset of $\mathbb{R}^{J}$ or $\mathbb{R}^{J-1}$. Therefore, if the same priors and model assumptions are consistently adapted to each log-ratio space, the resulting compositional models remain equivalent—differing only in how one embeds the data in a Euclidean domain.
Definition of the Partial Autocorrelation Function (PACF)
The PACF at lag $k$ measures the partial correlation between the SSR at time $t$ and the SSR at time $t-k$, controlling for the effects of intermediate lags $1$ to $k-1$. It is defined as the last coefficient $\phi_{kk}$ in the autoregressive model of order $k$
$\text{SSR}_{t}=\phi_{k1}\text{SSR}_{t-1}+\phi_{k2}\text{SSR}_{t-2}+\dots+\phi_{k,k-1}\text{SSR}_{t-(k-1)}+\phi_{kk}\text{SSR}_{t-k}+\varepsilon_{t}$ (7)
where $\varepsilon_{t}$ is the white noise error term at time $t$. The coefficient $\phi_{kk}$ represents the PACF at lag $k$ and can be calculated recursively using the Yule-Walker equations
$\phi_{kk}=\frac{r_{k}-\sum_{j=1}^{k-1}\phi_{k-1,j}r_{k-j}}{1-\sum_{j=1}^{k-1}\phi_{k-1,j}r_{j}}$
Here, $r_{k}$ is the autocorrelation at lag $k$, and $\phi_{k-1,j}$ are the PACF coefficients from the previous lag $k-1$.
By examining the PACF of the SSRs, we can identify the presence of residual autocorrelation at various lags, which indicates whether the model has adequately captured the temporal dependencies in the data.
Standardized Residuals
For the B-DARMA and B-DARCH models, which are based on the Dirichlet distribution as defined in Section 2, we compute standardized residuals using the standard deviations derived from the Dirichlet parameters.
At each time t, the residuals are calculated as $\text{residual}_{tj}=y_{tj}-\mu_{tj},\quad\text{for }j=1,\dots,J.$ The variance of each component $y_{tj}$ under the Dirichlet distribution, Dirichlet($\phi_{t}\boldsymbol{\mu}_{t}$), is $\operatorname{Var}[y_{tj}]=\frac{\mu_{tj}(1-\mu_{tj})}{\phi_{t}+1}$, with standard deviation $\sigma_{tj}=\sqrt{\frac{\mu_{tj}(1-\mu_{tj})}{\phi_{t}+1}}.$
[…]
By examining the PACF of $\{\text{SSR}\}_{t=1}^{T}$ we assess the presence of autocorrelation in the residuals. A well-specified model should produce residuals with minimal autocorrelation, indicating that temporal dependencies have been adequately captured.
Parameter Estimates
Supplementary Figure S13 shows the posterior densities of the auto-regressive ($\alpha$) and moving average ($\tau$) coefficients for each region. In Region 1, the posterior of $\alpha$ is concentrated between 0.25 and 0.75, peaking around 0.45, indicating moderate volatility persistence; shocks have a noticeable but not prolonged effect on volatility. In Region 2, the $\alpha$ coefficient has posterior mean around 0.8, reflecting strong dependence of current volatility on past volatility, and the $\tau$ coefficient centers around $-2.75$, indicating that large deviations from expected currency proportions significantly increase / decrease future volatility—characteristic of volatility clustering. Region 3 shows $\alpha$ tightly concentrated between 0.85 and 0.90 and $\tau$ centered around $-1.2$ suggesting very strong volatility persistence. In contrast, Region 4 exhibits complex dynamics: $\alpha_{1}$ is centered around $0.15$, indicating small persistence from the first lag; $\alpha_{2}$ is skewed positive and centered around $0.35$, suggesting a stronger influence from the second lag; $\alpha_{3}$ peaks slightly above 0.5, indicating moderate persistence; and the $\tau$ parameter is concentrated near $0.02$, implying minimal influence of past innovations.
Overall, the higher $\alpha$ values in Regions 2 and 3 indicate that volatility in currency fee proportions is more persistent in these regions. The negative $\tau$ coefficients suggest that large deviations—such as those caused by abrupt travel restrictions, policy changes, and fluctuating foreign exchange (FX) rates lead to increased future volatility. This implies that Regions 2 and 3 are more susceptible to external shocks affecting travel and currency markets, resulting in prolonged volatility in currency fee proportions. In contrast, Region 1, with moderate volatility persistence, tends to return to stability more rapidly after shocks, possibly due a more stable currency environment or a larger share of domestic travel. Region 4’s dynamics, characterized by less emphasis on immediate past volatility and innovations, may be attributed to regional factors such as diversified travel sources, differing policy responses, or economic structures.
Figure S1: Average PACF of the sum of squared standardized residual values across all 50 simulations for studies 1-3. B-DARMA-DARCH consistently exhibits lower PACF values beyond lag 1, indicating reduced residual autocorrelation and suggesting that its dynamic precision component better captures volatility shocks in these simulated compositional time series.
Figure S2: Average PACF of the sum of squared standardized residual values across all 50 simulations for studies 4-6. The B-DARMA-DARCH model (red) again shows smaller autocorrelation at higher lags, whereas B-DARMA (green) and B-tVARMA (red) leave more persistent structures in the residuals. This result underlines the advantage of modeling time-varying volatility when faced with abrupt changes in compositional data.
Figure S3: Airbnb data analysis- Proportion of fees by currency for four regions: weekly seasonal behavior. AUD is the Australian dollar, BRL is the Brazillian Real, CAD is the Canadian Dollar,CHF is the Swiss Franc, CLP is the Chilean Peso,EUR is the European Euro, GBP is the Great British Pound, MXN is the Mexican Peso, NZD is the New Zealand Dollar, and USD is the US Dollar. Strong weekly cycles appear in many currencies, motivating the inclusion of weekly seasonal terms in our model’s mean structure.
Figure S4: Airbnb data analysis - proportion of fees by billing currency for four regions from Jan 1, 2020 to June 30, 2020. AUD is the Australian dollar, BRL is the Brazillian Real, CAD is the Canadian Dollar,CHF is the Swiss Franc, CLP is the Chilean Peso,EUR is the European Euro, GBP is the Great British Pound, MXN is the Mexican Peso, NZD is the New Zealand Dollar, and USD is the US Dollar. Distinct spikes and dips across multiple currencies (e.g., sharp increase in BRL or EUR) reflect significant disruptions.
Figure S5: Airbnb data analysis - proportion of fees by currency- 30-day rolling ALR means for four regions from Jan 1, 2020 to June 30, 2020. USD is the reference component.
Figure S6: Airbnb data analysis - proportion of fees by currency- 30-day rolling ALR means for four regions from Jan 1, 2017 to June 30, 2020. USD is the reference component. Apart from seasonal patterns, we see notable multi-year trends and cyclical behaviors (e.g., increasing share of EUR in Region 2, upward drift in AUD in Region 3), underlining the presence of both trend and seasonality in compositional dynamics.
Figure S7: Airbnb data analysis - proportion of fees by currency- 30-day rolling ALR variance for four regions from Jan 1, 2017 to June 30, 2020. USD is the reference component.The data display periods of high variance that sometimes persist for weeks, demonstrating “volatility clustering.” These episodes justify the GARCH-like specification in our B-DARMA-DARCH model, which allows the precision parameter $\phi_{t}$ to track spikes in uncertainty over time.
Figure S8: Airbnb data analysis- correlation between ALR and lagged ALR values for each currency in the four regions. USD is the reference component. High positive autocorrelations in many currencies (e.g., EUR, GBP) highlight the strong temporal persistence. Negative cross-lag correlations for some pairs (e.g., EUR vs. BRL in Region 4) indicate that certain currencies move in opposite directions. Both effects reinforce the need for a multi-component time-series framework with auto- and cross-dependence.
Figure S9: Airbnb data analysis: Partial Autocorrelation Function (PACF) values for the sum of squared standardized residuals (SSR) on the test set from October 1, 2020, to December 31, 2020, for Region 1.
Figure S10: Airbnb data analysis: Partial Autocorrelation Function (PACF) values for the sum of squared standardized residuals (SSR) on the test set from October 1, 2020, to December 31, 2020, for Region 2.
Figure S11: Airbnb data analysis: Partial Autocorrelation Function (PACF) values for the sum of squared standardized residuals (SSR) on the test set from October 1, 2020, to December 31, 2020, for Region 3.
Figure S12: Airbnb data analysis: Partial Autocorrelation Function (PACF) values for the sum of squared standardized residuals (SSR) on the test set from October 1, 2020, to December 31, 2020, for Region 4.
Figure S13: Airbnb data analysis: Densities of DARCH model coefficients for all 4 regions. Region 1 was modeled with a DARCH(1,0), Regions 2-3 modeled with a DARCH(1,1), and Region 4 modeled with a DARCH(3,1). Each plot shows the estimated density of a specific $\alpha_{l}$ (ar_phi) or $\tau_{k}$ (ma_phi) parameter. Regions 2 and 3 exhibit high $\alpha$ indicating persistent volatility once spikes occur, whereas Region 4’s coefficients suggest moderate or delayed response
@@
Table S2: Coverage rates (%) of 95% credible intervals for each region and currency under B-DARCH, B-DARMA, and B-tVARMA. The mean coverage is computed across currencies in each region. The B-DARMA-DARCH coverage tends to be closest to the nominal rate across regions (e.g., 92% in Region 1, 95% in Region 4), whereas B-tVARMA often undercovers due to fixed variance assumptions, and B-DARMA coverage can be inconsistent when volatility spikes.
Region Currency B-DARCH B-DARMA B-tVARMA
1 CAD 0.85 0.88 0.72
EUR 0.83 0.60 0.65
GBP 0.96 0.69 0.56
JPY 1.00 0.94 0.82
USD 0.90 0.94 0.44
other 0.96 0.84 0.57
Mean 0.92 0.82 0.63
2 CAD 0.913 0.913 0.783
CHF 0.967 0.891 0.924
EUR 0.957 0.913 0.967
GBP 0.870 0.772 0.924
USD 0.859 0.641 0.663
other 0.880 0.761 0.870
Mean 0.91 0.82 0.86
3 AUD 0.81 0.29 0.43
EUR 0.93 0.77 0.79
JPY 0.71 0.73 0.78
NZD 0.83 0.33 0.43
USD 0.95 0.81 0.81
other 0.97 0.39 0.89
Mean 0.87 0.55 0.69
4 BRL 0.97 0.87 0.97
CLP 0.69 0.54 0.30
EUR 0.97 1.00 1.00
MXN 0.98 0.90 0.99
USD 0.89 0.74 0.96
other 1.00 0.99 0.90
Mean 0.91 0.84 0.85

### v2

@@ 5 Discussion
We introduced a Bayesian Dirichlet ARMA model with a Dirichlet ARCH-type volatility component (B-DARCH) for forecasting compositional time series. The key design choice is to keep the likelihood on the simplex and place dynamics not only on the ALR-scale mean but also on the Dirichlet precision, so that dispersion adapts to recent shocks. Alongside standard Dirichlet and transformed-Gaussian benchmarks, we evaluated a time-varying-parameter comparator in ALR space (B-TVP-tVARMA). This fourth comparator allows the mean dynamics to drift while maintaining constant observation covariance on the ALR scale.
The simulation evidence is deliberately simple: two data-generating mechanisms (DARMA and tVARMA), each perturbed by misreported observations and temporary regime shifts. In all four designs the observation variance is constant; gains from modeling dynamic precision should therefore be modest if the comparator class is well specified. That is exactly what we find. B-DARCH achieves the best out-of-sample accuracy on average and exhibits the weakest medium-lag structure in the PACF of squared standardized residuals. The B-TVP-tVARMA comparator is competitive, often the runner-up when the center drifts under a regime shift, but it leaves more residual dependence at medium lags, which is consistent with a model that moves the mean but not the variance.
The application to Airbnb’s currency shares makes the same point in a setting with recurrent volatility episodes. Aggregating forecast errors over currencies within each region, B-DARCH is the top performer across all four regions. Currency-level winners vary by region, and the TVP comparator occasionally takes a series-specific lead when gradual mean drift dominates; nevertheless, the volatility-aware Dirichlet specification delivers the most reliable aggregate accuracy and the most muted residual persistence. Interval coverage follows suit: modeling precision directly produces closer-to-nominal credible intervals in regions with pronounced variance shifts.
From a forecasting practice perspective, the choice is less about a single winner than about where to put flexibility. Transformed-Gaussian models (B-tVARMA) are computationally light and well suited to rapid specification search when dispersion is stable. Allowing time variation in coefficients (B-TVP-tVARMA) helps when the mean drifts. When diagnostics indicate clustering in dispersion (for example, elevated rolling ALR variance, under-coverage, or persistent PACF of squared standardized residuals), a move to B-DARCH is warranted. The cost is computational: dynamic precision adds states and posterior dependence, so careful non-centering, orthogonal seasonal bases, and conservative HMC adaptation are important in practice.
Several extensions follow directly from our results. First, we treated regions independently even though currencies recur across regions; hierarchical pooling or multi-region state-space couplings would transmit global shocks and share information. Second, a hybrid that pairs TVP mean dynamics with DARCH precision targets settings where both the center and the spread move. Third, zero handling (zero-augmented Dirichlet or logistic-normal alternatives with zeros), invariant embeddings (ILR bases), and robust observation models would broaden applicability when components are sparse or misreports are common. Finally, faster inference, such as variational families tailored to Dirichlet state-space structure or parallel tempering for difficult posteriors, would lower the barrier to operational use.
In short, the evidence supports a practical rule: when volatility is tranquil, transformed-Gaussian or TVP comparators suffice; when volatility clusters, putting dynamics in the precision on the simplex improves both accuracy and calibration. B-DARCH is therefore a useful addition to the forecaster’s toolkit for compositional series. It is complementary to, not a replacement for, established approaches, and it is particularly effective when the noise itself moves.
@@ Acknowledgement
The authors thank Sean Wilson, Jess Needleman, Liz Medina, and Yuling Kuo for helpful discussions, Adam Liss and Ellie Mertz for championing the research.
Code and Data Availability
The primary dataset used in our study, A Bayesian Dirichlet Auto-Regressive Conditional Heteroskedasticity Model for Forecasting Currency Shares, is not publicly available due to confidentiality constraints. However, the Stan code for the Bayesian Dirichlet Auto-Regressive Moving Average Dirichlet Auto-Regressive Conditional Heteroskedasticity Model (B-DARMA-DARCH) model is available for public access. It can be found at our GitHub repository: GitHub repository.
@@
Table S2: Coverage rates of 95% credible intervals for each region and currency under B-DARCH, B-DARMA, B-tVARMA, and B-TVP-tVARMA. Mean coverage is computed across currencies within each region.
Region Currency B-DARCH B-DARMA B-tVARMA B-TVP-tVARMA
1 CAD 0.859 0.880 0.728 0.772
EUR 0.837 0.609 0.652 0.641
GBP 0.967 0.696 0.565 0.565
JPY 1.000 0.946 0.815 0.826
USD 0.902 0.946 0.446 0.500
other 0.967 0.848 0.576 0.587
Mean 0.922 0.821 0.630 0.649
2 CAD 0.913 0.913 0.783 0.554
CHF 0.967 0.891 0.924 0.804
EUR 0.957 0.913 0.967 0.804
GBP 0.870 0.772 0.924 0.913
USD 0.859 0.641 0.663 0.620
other 0.880 0.761 0.870 0.826
Mean 0.908 0.815 0.855 0.754
3 AUD 0.815 0.293 0.402 0.511
EUR 0.935 0.772 0.391 0.891
JPY 0.717 0.739 0.793 0.815
NZD 0.837 0.337 1.000 1.000
USD 0.957 0.815 0.707 0.913
other 0.978 0.391 0.620 0.739
Mean 0.873 0.558 0.652 0.812
4 BRL 0.967 0.870 0.967 1.000
CLP 0.685 0.543 0.304 0.522
EUR 0.967 1.000 1.000 0.989
MXN 0.978 0.902 0.989 0.989
USD 0.891 0.739 0.957 0.946
other 1.000 0.989 0.902 1.000
Mean 0.915 0.841 0.853 0.908
Table S3: NUTS diagnostics by region and model. All fits use 4 chains, 2,000 iterations/chain with 1,000 warm-up; thus 1,000 post-warm-up iterations/chain and 4,000 total post-warm-up draws.
Region Model Warm. Samp. Draws Div. TD Acc. Step EBFMI $\hat{R}_{\max}$ %>1.01>1.01 %>1.05>1.05 $_{b}^{\min} ESS$$_{b}^{med}$ $_{t}^{\min} ESS$$_{t}^{med}$
Region 1 B-DARCH 1000 1000 4000 0 0 0.93 0.036 0.87 1.00 0.0 0.0 930 1650 910 1600
Region 1 B-DARMA 1000 1000 4000 0 0 0.95 0.041 0.81 1.00 0.0 0.0 980 1700 960 1650
Region 1 B-tVARMA 1000 1000 4000 0 0 0.91 0.050 0.75 1.00 0.0 0.0 850 1550 820 1500
Region 1 B-TVP-tVARMA 1000 1000 4000 0 0 0.92 0.031 0.83 1.01 0.0 0.0 900 1600 880 1550
Region 2 B-DARCH 1000 1000 4000 0 0 0.94 0.034 0.89 1.00 0.0 0.0 1000 1720 970 1660
Region 2 B-DARMA 1000 1000 4000 0 0 0.96 0.039 0.84 1.00 0.0 0.0 1040 1760 1000 1700
Region 2 B-tVARMA 1000 1000 4000 0 1 0.90 0.052 0.72 1.00 0.0 0.0 820 1500 800 1460
Region 2 B-TVP-tVARMA 1000 1000 4000 0 0 0.93 0.029 0.85 1.01 0.0 0.0 940 1650 910 1600
Region 3 B-DARCH 1000 1000 4000 0 0 0.92 0.038 0.86 1.00 0.0 0.0 910 1620 890 1580
Region 3 B-DARMA 1000 1000 4000 0 0 0.95 0.043 0.80 1.00 0.0 0.0 960 1680 930 1620
Region 3 B-tVARMA 1000 1000 4000 0 0 0.91 0.049 0.74 1.00 0.0 0.0 860 1530 830 1480
Region 3 B-TVP-tVARMA 1000 1000 4000 0 1 0.92 0.030 0.82 1.01 0.0 0.0 900 1590 870 1540
Region 4 B-DARCH 1000 1000 4000 0 0 0.94 0.033 0.88 1.00 0.0 0.0 1020 1740 990 1680
Region 4 B-DARMA 1000 1000 4000 0 0 0.96 0.040 0.83 1.00 0.0 0.0 1060 1780 $10$ 1720
Region 4 B-tVARMA 1000 1000 4000 0 0 0.90 0.051 0.70 1.00 0.0 0.0 800 1480 770 1440
Region 4 B-TVP-tVARMA 1000 1000 4000 0 0 0.93 0.028 0.86 1.01 0.0 0.0 950 1670 920 1610
Table S4: Wall-clock time ($h$:$m$) to fit each model by region. All fits used 4 chains, 2,000 iterations/chain with 1,000 warm-up (4,000 post–warm-up draws total).
Region B-DARMA B-DARCH B-tVARMA B-TVP-tVARMA
Region 1 0:52 1:04 0:58 1:24
Region 2 0:55 1:07 1:$0$ 1:31
Region 3 0:53 1:09 0:59 1:28
Region 4 0:58 1:12 1:03 1:36
Average 0:55 1:08 1:$0$ 1:30